# Date of publication xxxx 00, 0000, date of current version xxxx 00, 0000. PREPRINT of Accepted Manusript

### v4

@@ III-E Implementation Details
For the contrastive SSL pretraining phase, we train the conv layers of the network by projecting the flattened output of the last layer, with an original dimension of into a 128-D space via a dense or fully-connected layer and by ultimately minimizing the NT-Xent loss (eq. 1). The dimension of the projection was set to follow the one in the original SimCLR paper [70]. For the cosine similarity (eq. 2) we use a temperature of 0.1, as in [27], while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer [73] with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.

### v5

@@ III-E Implementation Details
For the contrastive SSL pretraining phase, we train the convolutional layers of the network by projecting the flattened output of the last layer, with an original dimension of into a 128-D space via a dense or fully-connected layer and by ultimately minimizing the NT-Xent loss (eq. 1). The dimension of the projection was set to follow the one in the original SimCLR paper [70]. For the cosine similarity (eq. 2) we use a temperature of 0.1, as in [27], while the initial batch size is set to 256 (leading to 512 after augmentation occurs). The SSL training epochs are set to a maximum of 200 epochs, based on an early stopping set on the validation split with a patience of 10 epochs. Additionally, to accommodate the large batch size we employ the LARS optimizer [73] with a linear learning rate warm-up of 20 epochs, up to 0.1, followed by a cosine decay of 1%.